Driving at 35 mph, one covers 18 miles in how many seconds?
Answer by stanbon(75887)

Driving at 35 mph, one covers 18 miles in how many seconds?
----
time = distance/rate = 18/35 hrs = (18/35)(3600 seconds) = 1851.43 seconds
=====================
Cheers,
Stan H.